A man is 40 years old and his son is 8 years old. in hiw many years will the man be 3 times as old as his son will be then?
Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

x years away
Man will be 40+x, son 8+x
The man is 3 times the age of his son then, so that 40+x=3(8+x)=3x+24
2x=16
x=8 years
man will be 48 and son 16.

Answer by greenestamps(13206)



(1) The difference in their ages is 40-8 = 32 years.

(2) When the man is 3 times as old as his son, the difference in their ages will be 3x-x=2x.

So

2x+=+32
x+=+16

The man will be 3 times as old as his son when his son is 16.

Since his son is now 8, that will be 16-8 = 8 years from now.